OpenStudy (yacoub1993):

21. Write an equation of a line whose graph is parallel to the graph of y = 3x – 10.

OpenStudy (anonymous):

well,you know when a line is parallel to another line , their slopes are equal to each other?

OpenStudy (yacoub1993):

yes

OpenStudy (anonymous):

ok,so the slope of the first one is "3" so the equation is: y = 3x + b

OpenStudy (anonymous):

but we don't have b, okay?

OpenStudy (yacoub1993):

so my answer of y=3x + 5 would be correct for the question?

OpenStudy (anonymous):

why 5?

OpenStudy (anonymous):

how did you find b=5?

OpenStudy (yacoub1993):

i am assuming right now not sure

OpenStudy (anonymous):

ok,the question just gave you slope,and didn't give you any point,so we can't find y-intercept...

OpenStudy (yacoub1993):

then

OpenStudy (yacoub1993):

@Yttrium how do i solve this one

OpenStudy (anonymous):

you mean another question?

OpenStudy (anonymous):

multiply the equation with any constant and you'll get a parallel line.

OpenStudy (yttrium):

You already know the slope, right? Therefore, you can substitute any point that do not touch the given equation.

OpenStudy (yacoub1993):

example ?@Yttrium

OpenStudy (yttrium):

(0,0) then substitute it to the form y-y1 = m (x-x1) where your m=3

OpenStudy (anonymous):

parallel lines have equal slopes, this means that we know that the slope of the unknown parallel line is..? :D

OpenStudy (anonymous):

If you "don't use formula" it is better,or if you want to use formula "you need to know its proof"

OpenStudy (anonymous):

please don't use any formula without any knowledge of its proof :)

OpenStudy (anonymous):

for example you can solve @Yttrium 's question like this: y = m(x) + b and you know the slope,y and x so you can find b simply :)

OpenStudy (yacoub1993):

ok wait let me solve it

OpenStudy (yacoub1993):

0 = 3(0) + b 0 = 0 + b 0 = b

OpenStudy (yacoub1993):

is it like that

OpenStudy (anonymous):

correct :)

OpenStudy (yacoub1993):

let me try my equation if i get it right

OpenStudy (yacoub1993):

nah failed could not get any

OpenStudy (yacoub1993):

@Yttrium give me 1 so that i answer it there

OpenStudy (yttrium):

@Yacoub1993, you already had an answer. Since you've got b=0. Therefore, y=3x+ or simply y=3x is another equation with slope of 3. :)

OpenStudy (yacoub1993):

ok thnxs